To determine the enthalpy change for a metal displacement reaction by adding a known mass of zinc powder to excess aqueous copper(II) sulfate solution and recording the rise in temperature.
Initial mass of weighing boat / g = 2.54
Mass of weighing boat and zinc / g = 4.57
Final mass of weighing boat / g = 2.55
Mass of zinc added / g = 2.02
Volume of aqueous CuSO4 used = 40 cm3
Temperature change / ºC = 29
1
Expert's answer
2020-10-26T14:57:38-0400

To determine the enthalpy change for a metal displacement reaction by adding a known mass of zinc powder to excess aqueous copper(II) sulfate solution and recording the rise in temperature.

Initial mass of weighing boat / g = 2.54

Mass of weighing boat and zinc / g = 4.57

Final mass of weighing boat / g = 2.55

Mass of zinc added / g = 2.02

Volume of aqueous CuSO4 used = 40 cm3

Temperature change / ºC = 290C


Solution:


mass of zinc reacted = 2.02g.

Volume of aqueous CuSO4 used = 40cm3 .

Temperature change (0C) = 290C.


The reaction of Zn metal and CuSO4 (aq) is given as


Zn(s) + CuSO4 (aq) ---> ZnSO4 (aq) + Cu(s)


Our first step will be to find the Heat produced due to displacement of copper by zinc.


Step 1: Find the heat produced due to the displacement of copper by zinc.


The formula which relates heat and specific heat capacity is


Q = m * s * ΔT; ------------------------------- Equation 1


where,

m = mass of water = volume of copper sulphate solution (cm3) = 40cm3.

s = specific heat capacity of the solution (water) = 4.184 J/mol-K.

ΔT = change in temperature = 290C.


The density of water is 1.00g/cm3

So mass of water = 40cm3 * 1.00g/cm3 = 40g


Substitute, m= 40g, s = 4.184 J/g C, and ΔT = 29 0C in the Equation 1, we have


Q = 40g * 4.184J/g0C * 290C


Q = 4853.44J.


Step 2: To find the moles of zinc reacted in the reaction.


In the given experiment known amount of Zinc is used and CuSO4 solution used is in excess.

So Zinc is the limiting reactant and we will use its moles for finding the molar enthalpy of reaction.


Atomic mass of zinc = 65.38g/mol


moles of zinc = 2.02g of Zn * 1mol Zn/ 65.38g Zn = 0.03090 mol Zn.



Step 3: To find the molar enthalpy of the reaction.


Molar enthalpy of displacement of copper by zinc = -Q/ (mol of Zn)

= - 4853.44J / 0.03090 mol Zn

= - 157088 J/mole


Convert this to ‘kJ/mol’ using 1kJ = 1000J.


Molar enthalpy (kJ/mol) = - 157088 J/1mole * 1kJ/1000J = -157.088 kJ/mol

Hence molar enthalpy for the displacement of copper by zinc is -157 kJ/mol